HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

SUB-MILLIMETER WAVE SPECTROSCOPY OF THE CH_{2}CHO RADICAL

\underline{R. CHAHBAZIAN} , \underline{M.-A. MARTIN-DRUMEL} , \underline{O. PIRALI} , Université Paris-Saclay, CNRS, Institut des Sciences Moléculaires d’Orsay, 91400 Orsay,

France

The vinoxy radical (CH_{2}CHO) is a known product of the reaction involving O(^{3}P) and unsaturated hydrocarbons as C_{2}H_{4}^{1}. Its detection in the interstellar medium would help to understand the chemical networks involving large radicals.

The CH_{2}CHO radical has previously been studied at high resolution by several authors. Its pure rotational spectrum was first investigated by Endo and Nakajima^{2} using millimeter-wave spectroscopy and later by Endo, Saito, and Hirota^{3} using Fourier-transform microwave spectroscopy and double resonance spectroscopy. Utkin et al.^{4} studied its rovibrational spectra allowing for ground state combination differences.

In this work, we have measured the pure rotational spectrum of the vinoxy radical at millimeter- and sub-millimeter wavelengths (110–860 GHz). CH_{2}CHO was produced by H abstraction from acetaldehyde (CH_{3}CHO) using atomic fluorine in a double pass absorption cell at room temperature^{5}. A Zeeman-modulated spectrometer has been used to record the pure rotational transitions of the radical, in which a magnetic field is modulated in magnitude. As the recorded spectra are free from signals arising from closed-shell species, the setup allows spectral acquisition over wide spectral windows. We probed transitions involving relatively high values of the rotational quantum numbers N′′ and K′′_{a} (up to 41 and 18 respectively). Their combination with available high-resolution literature data (pure rotation^{1,2} and ground state combination difference from ro-vibration studies^{3}) led to an improvement of the modeling of the CH_{2}CHO rotational structure. This laboratory study provides complete spectroscopic information to search for CH_{2}CHO in various interstellar environments, from cold (e.g. typically 10 K for dense molecular clouds) to warm (e.g., ∼ 200 K for hot corinos) objects.

1A. Occhiogrosso, S. Viti, and N. Balucani, Monthly Notices of the Royal Astronomical Society, 432 3423 (2013)

2Y. Endo and M. Nakajima, Journal of Molecular Spectroscopy 301 15 (2014)

3Y. Endo, S. Saito, and E. Hirota, The Journal of Chemical Physics 83 2026 (1985)

4Y. G. Utkin, J.-X. Han, F. Sun, H.-B. Chen, G. Scott, and R. F. Curl, The Journal of Chemical Physics, 118 10470 (2003)

5G. Inoue and H. Akimoto, J. Chem. Phys. 74 425 (1981)

p-number: p051  Submitted on Thu, 08 Jun 2023 14:44:58 +0200